### v1

# TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning

Abstract
Retrieval-Augmented Generation (RAG) has demonstrated considerable effectiveness in open-domain question answering. However, when applied to heterogeneous documents, comprising both textual and tabular components, existing RAG approaches exhibit critical limitations. The prevailing practice of flattening tables and chunking strategies disrupts the intrinsic tabular structure, leads to information loss, and undermines the reasoning capabilities of LLMs in multi-hop, global queries. To address these challenges, we propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data. TableRAG iteratively operates in four steps: context-sensitive query decomposition, text retrieval, SQL programming and execution, and compositional intermediate answer generation. We also develop HeteQA, a novel benchmark designed to evaluate the multi-hop heterogeneous reasoning capabilities. Experimental results demonstrate that TableRAG consistently outperforms existing baselines on both public datasets and our HeteQA, establishing a new state-of-the-art for heterogeneous document question answering. We release TableRAG at https://github.com/yxh-y/TableRAG/tree/main.
TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning
Xiaohan Yu∗ Huawei Cloud BU, Beijing yuxiaohan5@huawei.com Pu Jian∗ Huawei Cloud BU, Beijing jianpu2@huawei.com Chong Chen† Huawei Cloud BU, Beijing chenchong55@huawei.com
1 Introduction
Heterogeneous document-based question answering (Chen et al., 2020), which necessitates reasoning over both unstructured text and structured tabular data, presents substantial challenges. Tables are characterized by interdependent rows and columns, while natural language texts are sequential. Bridging this divergence within a unified QA system remains a non-trivial task.
The prevailing approach extends the retrieval-augmented generation (RAG) paradigm, in which the tables are linearized into textual representations (e.g., Markdown) (Gao et al., 2023; Jin and Lu, 2023; Ye et al., 2023b). Typically, chunking strategies are employed (Finardi et al., 2024), wherein flattened tables are segmented and merged with adjacent text spans. During inference, the LLMs generate answers based on the top-N retrieved chunks. However, these methodologies are predominantly tailored to scenarios that require only surface-level comprehension of tables, such as direct answer extraction Pasupat and Liang (2015a); Zhu et al. (2021). When applied to extensive documents that interleave textual and tabular elements, existing RAG methodologies exhibit critical limitations:
Structural Information Loss: The tabular structure integrity is compromised, leading to information loss or irrelevant context that impedes downstream LLMs performance.
Lack of Global View: Due to document fragmentation, the RAG system struggles with multi-hop global queries Edge et al. (2024), such as aggregation, mathematical computations, and other reasoning tasks that require a holistic understanding across entire tables.
As illustrated in Figure 1, the RAG approach computes percentage over the top-N most relevant chunks rather than the full table, and thus results in an incorrect answer.
To address these limitations of existing RAG systems, we propose TableRAG, an hybrid (SQL execution and text retrieval) framework that dynamic transitions between textual understanding and complex manipulations over tabular data. TableRAG interacts with tables by leveraging SQL as an interface. Concretely, the framework operates via a two-stage process: an offline database construction phase and an online inference phase of iterative reasoning. The iterative reasoning procedure comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) SQL programming and execution, and (iv) intermediate answer generation. The utilization of SQL enables precise symbolic execution by treating table-related queries as indivisible reasoning units, thereby enhancing both computational efficiency and reasoning fidelity. To facilitate rigorous evaluation of multi-hop reasoning over heterogeneous documents, we introduce HeteQA, a novel benchmark consisting of 304 examples across nine diverse domains. Each example contains a composition across five distinct tabular operations. We evaluate TableRAG on both established public benchmarks and our HeteQA dataset against strong baselines, including generic RAG and program-aided approaches. Experimental results demonstrate that TableRAG consistently achieves state-of-the-art performance. Overall, our contributions are summarized as follows:
We identify two key limitations of existing RAG approaches in the context of heterogeneous document question answering: structural information loss and lack of global view.
We propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data. TableRAG comprises an offline database construction phase and a four-step online iterative reasoning process.
We develop HeteQA, a benchmark for evaluating multi-hop heterogeneous reasoning capabilities. Experimental results show that TableRAG outperforms RAG and programmatic approaches on HeteQA and public benchmarks, establishing a state-of-the-art solution.
2 Task Formulation
In the context of the heterogeneous document question answering task, we define the task input as extensive documents, denoted as where denotes the textual contents and refers to the tabular components. Given a user question , the objective of this task is to optimize a function that, given the combined textual and tabular context, can produce the correct answer :
3 TableRAG Framework
3.1 Overview and Design Principles
We propose TableRAG, an hybrid (SQL execution and text retrieval) framework designed to preserve table structural integrity and facilitate heterogeneous reasoning. As depicted in Figure 2, TableRAG consists of offline and online workflows. The offline phase is tasked with database construction, while the online phase facilitates iterative reasoning. The reasoning procedure unfolds in a four-stage process: (i) Context-sensitive query decomposition, which identifies the respective roles of textual and tabular modalities within the query. (ii) Text retrieval. (iii) SQL programming and execution, which is selectively invoked for subqueries requiring tabular data reasoning. (iv) Compositional intermediate answer generation. The preferential use of SQL is motivated by its capacity to leverage the expressive strength of symbolic execution over structured data, thereby enabling tabular components within user queries to be treated as monolithic reasoning units. In contrast, other languages like Python incur substantial computational overhead when dealing with large-scale data or complex workloads (Shahrokhi et al., 2024).
3.2 Database Construction
In the offline stage, we first extract structured components from heterogeneous documents, yielding a set of tables . To enable information retrieval, we construct two parallel corpora: a textual knowledge base and a tabular schema database. The textual knowledge base comprises both the raw texts and the Markdown-rendered form of each table, denoted as . Both and are segmented into chunks, which are then embedded into dense vector representations using a pre-trained language model (Chen et al., 2024a). For tabular schema database construction, we represent each table by a standardized schema description , derived via a template as follows:
Then, we define a mapping from each flattened table chunk to its originating table schema:
where denotes the -th chunk derived from table . This mapping ensures that local segments remain contextually anchored to the table structure from which they are derived.
The tables are also ingested in a relational database (e.g., MySQL111https://github.com/mysql), supporting symbolic query execution in the subsequent online reasoning.
3.3 Iterative Reasoning
To address multi-hop, global queries that require compositional reasoning over texts and tables, we introduce an iterative inference process aligned with in Equation 1. This process comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) program and execute SQL, and (iv) compositional intermediate answer generation. Through repeated cycles of decomposition and resolution, a solution to the query is progressively constructed. Detailed prompt templates are provided in Appendix F.
Context-Sensitive Query Decomposition
We explicitly delineate the respective roles of textual and tabular modalities during the reasoning process. While a table-related query may involve multiple semantic reasoning steps, its tabular resolution can collapse to a single executable operation. Consequently, an effective decomposition of global queries demands more than mere syntactic segmentation, but also structural awareness of the underlying data sources. To this end, we first retrieve the most relevant table content from the textual database and link it to its corresponding table schema description via the mapping function . Based on this, we formulate a subquery at the -th iteration.
Text Retrieval
We deploy a retrieval module that operates in two successive stages: vector-based recall followed by semantic reranking. Given an incoming query , it is encoded into a shared dense embedding space alongside document chunks. We then select the top- candidates with the highest cosine similarity to the query embedding:
In the subsequent reranking stage, the recalled candidate chunks are re-evaluated by a more expressive relevance model, yielding the final top- selections, denoted by .
SQL Programming and Execution
To support accurate reasoning over tabular data, we incorporate a "program-and-execute" mechanism that is selectively invoked only when subquery reasoning involves tables. Specifically, we inspect whether any content originates from tabular sources in the retrieved results. For each chunk in the top-ranked set , we apply the mapping function (in Equation 2) to extract its associated schema, yielding a table schema set:
If the set is empty, this module is passed. Otherwise, we derive an accurate answer with the current subquery and the corresponding schema context as inputs. To achieve this, we leverage structured query execution over relational data and use SQL as the intermediate formal language. A dedicated tool with LLM as backend generates executable SQL programs and applies them to the pre-constructed MySQL database, formalized as follows:
Intermediate Answer Generation
For the subquery , TableRAG can benefit from two heterogeneous information sources: the execution result over SQL database and text retrieval result from the document database. Both of the data sources provide partial or complete evidence. They introduce distinct failure modes: SQL execution may produce incorrect results or execution errors, while text retrieval may yield incomplete or misleading context. Consequently, the results from these sources may either reinforce each other or present contradictions. To address this, we adopt a compositional reasoning mechanism. The execution result and the retrieved textual chunks are cross-examined to validate consistency and guide answer selection. The final answer to each subquery is derived by adaptively weighting the reliability of each source based on its evidential utility, .
Once the query decomposition module determines that no further subqueries are necessary, TableRAG terminates the iterative reasoning process, yielding the final answer , where denotes the total number of iterations performed.
4 Benchmark Construction
In this section, we present HeteQA, a novel benchmark for assessing multi-hop reasoning across heterogeneous documents.
4.1 Data Collection
HeteQA necessitates advanced operations, such as arithmetic computation, nested logic, etc. To balance annotation fidelity with scalability, we adopt a human-in-the-loop collaborative strategy that integrates LLMs with human verification. The construction pipeline proceeds in three stages:
Query Generation
We curate tabular sources from the Wikipedia dataset (Chen et al., 2020). To facilitate analytical depth, we restrict our selection to tables with a minimum of 20 rows and 7 columns, and apply structural deduplication to eliminate redundancy across similar schemas. For each retained table, we define a suite of advanced operations, e.g., conditional filtering, and statistical aggregation. These operations serve as primitives for constructing complex queries. Leveraging the Claude-3.7-sonnet 222https://www.anthropic.com/claude/sonnet, we prompt for query synthesis as compositions over these primitives. Each generated query is paired with executable code in both SQL and Python. We execute the associated code and obtain the answer. A final deduplication pass is applied over both queries and answers, promoting diversity in the dataset. Full implementation details are provided in Appendix A.
Answer Verification
To ensure correctness and reliability, each instance is subjected to manual inspection by human annotators. Their task is to verify that the execution outcome is accurate for the corresponding query. In cases where discrepancies are found, they are responsible for correcting both the underlying code and the resulting answer.
Document Reference
To support queries that integrate both tabular and textual information, we augment the instance by leveraging the associated Wikipedia document. Specifically, certain entities within the query are replaced with reference-based formulations by the human annotators. For example, the query "Which driver " can be rephrased as "What is the nationality of the driver ". This entity substitution can either modify the subject of the question and its corresponding answer or alter the query phrasing while preserving the original answer. The annotation guidelines and annotator profiles are detailed in Appendix B.
4.2 Discussion
Each data instance in HeteQA is composed of a query, its corresponding answer, the executable SQL sentence, and the execution-derived answer. Through our data collection pipeline, we construct 304 high-quality examples whose answers are grounded in both single-source (82%) and multi-source (18%). The resulting benchmark spans 136 distinct tables and 5314 wiki knowledge entities. To characterize the dataset, we analyze its semantic domains and the types of tabular reasoning operations. As illustrated in Figure 3, HeteQA covers 9 semantically diverse domains and encompasses 5 principal categories of tabular operations. Together, HeteQA constitutes a structurally diverse and semantically broad resource for advancing question answering over heterogeneous documents.
5 Experiments
| Method | Backbone | HybridQA | WikiTQ | HeteQA | ||
| - | - | Single-Source | Multi-Source | Overall | ||
| Direct | Claude-3.5 | 9.84 | 6.21 | 10.68 | 8.65 | 10.00 |
| DeepSeek-R1 | 24.42 | 12.20 | 3.40 | 13.46 | 6.77 | |
| DeepSeek-V3 | 14.75 | 10.39 | 6.80 | 28.85 | 14.19 | |
| Qwen-2.5-72b | 11.47 | 7.37 | 4.85 | 12.50 | 7.42 | |
| NaiveRAG | Claude-3.5 | 20.28 | 82.60 | 33.20 | 40.35 | 34.54 |
| DeepSeek-V3 | 26.56 | 75.40 | 33.60 | 45.61 | 35.85 | |
| Qwen-2.5-72b | 22.62 | 66.33 | 23.07 | 36.84 | 25.66 | |
| ReAct | Claude-3.5 | 43.38 | 69.81 | 26.40 | 44.44 | 29.60 |
| DeepSeek-V3 | 38.36 | 63.40 | 21.14 | 47.39 | 26.07 | |
| Qwen-2.5-72b | 37.38 | 53.80 | 16.94 | 35.71 | 20.47 | |
| TableGPT2 | 9.51 | 63.40 | 35.60 | 16.67 | 32.24 | |
| TableRAG | Claude-3.5 | 47.87 | 84.62 | 44.94 | 40.74 | 44.19 |
| DeepSeek-V3 | 47.87 | 80.40 | 43.32 | 51.85 | 44.85 | |
| Qwen-2.5-72b | 48.52 | 78.00 | 37.65 | 43.96 | 38.82 | |
5.1 Experimental Settings
5.1.1 Datasets.
We assess the performance of TableRAG on our curated HeteQA, as well as multiple established benchmarks spanning two settings:
HybridQA
(Chen et al., 2020) A multi-hop QA dataset involving both tabular and textual information. For our evaluation, we only retain data cases with tables containing more than 100 cells.
WikiTableQuestion
(Pasupat and Liang, 2015b) A TableQA dataset spanning diverse domains. The queries necessitate a range of data manipulation operations, including comparison, aggregation, etc.
5.1.2 Implementation Details
In the text retrieval process, we employ the BGE-M3 series models (Chen et al., 2024a, b). During recall, we retain the top 30 candidates, from which the top 3 are subsequently selected via reranking. To manage large inputs, the text is chunked into segments of 1000 tokens, with a 200-token overlap between consecutive chunks. The iterative loop is bounded by a maximum of 5 iterations. For backbone LLMs, we utilize Claude-3.5-Sonnet as a representative closed-source LLM, while Deepseek-V3, Deepseek-R1 (Guo et al., 2025), and Qwen-2.5-72B (Yang et al., 2024) serve as the open-source counterparts. A consistent backend is maintained for all modules in the online iterative reasoning process. We use accuracy as the evaluation metric, assessed by Qwen-2.5-72B, which yields a binary score of 0 or 1. The prompt is shown in Appendix F.
5.1.3 Baselines
We evaluate the performance of TableRAG by benchmarking it against three distinct baseline methodologies: (1) Direct answer generation with LLMs. (2) NaiveRAG, which processes tabular data as linearized Markdown formatted texts and subsequently applies a standard RAG pipeline. (3) React (Yao et al., 2023), a prompt based paradigm to synergize reasoning and acting in LLMs with external knowledge sources. (4) TableGPT2 (Su et al., 2024) employs a Python-based execution module to generate code (e.g., Pandas) for answer derivation within a simulated environment. The detailed implementation of these baseline methods is provided in Appendix C.
5.2 Main Result
The main results across different LLMs backbones are presented in Table 1. Several key observations emerge: (1) The ReAct framework demonstrates advantages over naive RAG on multi-source data, but exhibits degraded performance on single-source data that requires tabular reasoning. This can be attributed to context insensitivity during multi-turn reasoning. Queries decomposed into multiple sub-tasks, such as filtering or aggregation, suffering from information incompleteness or error propagation. (2) TableGPT2 yields acceptable results solely on single-source queries, such as WikiTQ, underscoring its limited capacity for handling multi-source queries. This reflects a lack of generalizability in heterogeneous information environments. (3) TableRAG surpasses all baselines, achieving at least a 10% improvement over the strongest alternative. Notably, it performs robustly across both single-source and multi-source data. This performance gain is attributed to the incorporation of symbolic reasoning, which enables effective adaptation to heterogeneous documents. Moreover, the consistency in performance across different LLM backbones underscores TableRAG’s architectural generality and compatibility with a broad range of backbones.
5.3 Ablation Study
To elucidate the relative importance of each component within the TableRAG framework, we evaluate the full architecture against three ablated variants: (1) w/o Context-Sensitive Query Decomposition, where query decomposition is performed without conditioning on retrieved table schema. (2) w/o SQL Execution, which replaces the SQL programming and execution module with the markdown table format. (3) w/o Textual Retrieval, which operates solely through table-based SQL execution, without leveraging textual resources such as Wikipedia documents. The results are summarized in Figure 4. All the modules contribute to the overall performance of TableRAG, though their relative impact varies across benchmarks. On HybridQA, document retrieval proves particularly critical, due to its emphasis on the extraction of entity-centric or numerical cues. Conversely, for HeteQA, SQL execution proves more influential, as the queries involve nested operations that benefit from SQL-based symbolic reasoning. These findings highlight the complementary design of TableRAG’s textual retrieval and program-executed reasoning components.
6 Efficiency
We evaluate the efficiency of TableRAG by examining the distribution of its execution iterations, as illustrated in Figure 5. Execution lengths are grouped into four categories: fewer than 3 steps, 3–5 steps, exactly 5 steps, and more than 5 steps. Among the evaluated methods, TableGPT2 demonstrates the highest average number of execution steps, with a modal value centered around five. In contrast, TableRAG consistently requires fewer steps, resolving approximately 63.55% of instances in fewer than five steps and an additional 30.00% precisely within five, with only a marginal proportion of cases remaining unsolved under the given iteration constraints. While ReAct exhibits a comparable distribution in execution steps, its overall performance remains markedly inferior to that of TableRAG. These results suggest that TableRAG achieves both superior efficiency in execution and outstanding reasoning accuracy. It is attributed to the incorporation of SQL-based tabular reasoning.
7 Analysis
We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix D.
7.1 Error Analysis
In addition to evaluating the overall performance of TableRAG against established baselines, we performed a detailed error analysis to characterize the nature of prediction failures. Broadly, the incorrect outputs fall into two primary categories: (1) reasoning failures, attributable to errors in SQL execution or flawed intermediate query decomposition, and (2) task incompletion, typically manifesting as refusals to answer or termination upon exceeding the maximum iteration limit. The prediction distribution is shown in Figure 6. Notably, TableGPT2 exhibits the highest frequency of such failures, largely due to its limited capacity to integrate contextual cues from the wiki documents. This constraint frequently results in the model either explicitly refusing to respond or acknowledging its inability to do so. In contrast, ReAct, which lacks mechanisms for context-aware query decomposition and code execution simulation, often engages in unnecessarily elaborate reasoning steps for problems that could be addressed via a single structured inquiry. TableRAG demonstrates the lowest failure rate among the methodologies assessed. Its consistent ability to yield valid responses within five iterations highlights the efficacy of its design — particularly its use of context-aware query decomposition and selective SQL-based execution planning.
7.2 Prediction across Domains
Figure 7 presents a comparative evaluation of TableRAG, instantiated with various backbone LLMs, against the ReAct framework across various domains. The results reveal that TableRAG consistently outperforms ReAct in the majority of domains, demonstrating its effectiveness in heterogeneous document question answering. Only certain domains, such as Culture, exhibit comparatively weaker performance on TableRAG with Qwen backbone. A closer inspection of the data distribution suggests that this degradation may stem from the sparsity of domain-specific instances.
8 Related Work
8.1 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has emerged as a robust paradigm for mitigating hallucination (Zhang et al., 2023a) and enhancing the reliability of Large Language Models (LLMs) generated responses (Lewis et al., 2020; Guu et al., 2020). The RAG approaches retrieve from the knowledge base, and the most relevant document chunks are subsequently incorporated into the generation process (Gao et al., 2023; Zhu et al., 2024; Borgeaud et al., 2022). However, this straightforward retrieval process often yields noisy chunks that may lack critical details, thereby diminishing the quality of the subsequent generation. Recent advancements have thus focused on task-adaptive retrieval mechanisms. Notable frameworks in this regard include Self-RAG (Asai et al., 2023), RQ-RAG (Chan et al., 2024), etc. Despite these innovations, RAG still faces challenges when dealing with heterogeneous contexts (Satpute et al., 2024).
8.2 Table Reasoning via Large Language Models
Table reasoning refers to the development of a system that provides responses to user queries based on tabular data (Lu et al., 2025). The mainstream approaches to table reasoning can be broadly classified into two categories. The first category revolves around leveraging LLMs through prompt engineering. For instance, Tab-CoT (Jin and Lu, 2023) applies chain-of-thought (CoT) reasoning to establish a tabular structured reasoning process. Similarly, Chain-of-Table (Wang et al., 2024) extends the CoT methodology to the tabular setting, enabling a multi-step reasoning process for more complex table-based queries. The second category involves utilizing programs to process tabular data. Tabsqlify (Nahid and Rafiei, 2024) employs a text-to-SQL approach to decompose tables into smaller, contextually relevant sub-tables. DATER (Ye et al., 2023a) adopts a few-shot prompting strategy to reduce large tables into more manageable sub-tables, using a parsing-execution-filling technique that generates intermediate SQL queries. BINDER (Cheng et al., 2022; Zhang et al., 2023b) integrates both Python and SQL code to derive answers from tables. InfiAgent-DABench (Hu et al., 2024) utilizes an LLM-based agent that plans, writes code, interacts with a Python sandbox, and synthesizes results to solve table-based questions.
9 Conclusion
We address the limitations of existing RAG approaches in handling heterogeneous documents that combine textual and tabular data. Current approaches compromise the structural integrity of tables, resulting in information loss and degraded performance in global, multi-hop reasoning tasks. To overcome these issues, we introduce TableRAG, an SQL-driven framework that integrates textual understanding with precise tabular manipulation. To rigorously assess the capabilities of our approach, we also present a new benchmark HeteQA. Experimental evaluations across public datasets and HeteQA reveal that TableRAG significantly outperforms existing baseline approaches.
Limitations
While TableRAG demonstrates strong performance, several limitations merit consideration: 1. The effectiveness of TableRAG is closely tied to the capabilities of the underlying LLMs. Our implementation leverages high-capacity models such as Claude, DeepSeek-v3, and Qwen-72B-Instruct, which possess strong generalization abilities. Smaller models that lack specialized instruction tuning may exhibit a marked degradation in performance. This suggests that achieving competitive results may necessitate substantial computational resources. 2. The HeteQA benchmark is restricted to English. This limitation arises from the difficulty in curating high-quality heterogeneous sources across multiple languages. As a result, cross-lingual generalization remains unexplored. In future work, we aim to extend HeteQA to a multilingual setting, thereby broadening the applicability and robustness of our evaluation framework.
Appendix A HeteQA
A.1 Prompt Design
The templates employed for HeteQA benchmark construction are presented in Table LABEL:SQL.
A.2 Table Collection
To enable complex reasoning over tabular structures, we curate a subset of extensive tables from the Wikipedia-based corpus. Specifically, we retain only tables containing more than 20 rows and at least 7 columns. This filtering process reduced the initial collection from 15,314 tables to 1,345 candidates that meet the criteria for structural richness. To further enhance query diversity and reduce redundancy in the dataset, we apply a deduplication step based on schema similarity. In cases where multiple tables share identical column structures — for example, the entries for 1947 BAA draft and 1949 BAA draft — only a single representative instance is preserved. This procedure yielded a final dataset comprising 155 unique tables.
A.3 Data Collection
As illustrated in Figure 8, each data instance is constructed through a three-stage pipeline, facilitated via the Claude 3.7 Sonnet API333https://www.anthropic.com/claude/sonnet.
To construct the dataset, we prompt the LLM to generate SQL or pandas code associated with queries conditioned on a provided table schema. The generated code is then executed against the table to obtain the corresponding answer. Instances that fail to produce an executable result are discarded. To promote diversity within the dataset and avoid redundancy, we eliminate any instances exhibiting duplicate queries or identical answers. This filtering is performed using regular expressions to detect lexical or semantic repetition. Additionally, we discard cases where the execution result is ambiguous, such as queries seeking the top-ranked entity when multiple candidates are tied. Following automated filtering, all remaining instances undergo a manual verification process. Two human annotators independently assess each query, the corresponding code, and the resulting answer. In cases of inconsistency or error, annotators revise the SQL or Python code and correct the associated answer to ensure accuracy and coherence.
Appendix B Guidelines for HeteQA Annotations
B.1 Annotator Profiles
The two annotators are volunteers whose native language is Chinese and who possess fluent English proficiency. Both annotators are professional software engineers with substantial experience in SQL and Python programming.
B.2 Guidelines for Answer Verification
Objective
Your task is to verify whether the provided answer correctly and completely addresses the natural language query, based on the associated code and table. You will also identify and handle ambiguous or non-unique cases.
Annotation Steps
1. Verify the Answer
Read the natural language query and examine the accompanying Excel table, SQL statements, and Python (pandas) code. Verify whether the output aligns with the semantics of the original query and the underlying data. To this end, you are permitted to execute multiple operations on the provided Excel files.
2. Correction Tasks (if necessary)
If the answer is incorrect, modify both the code and the answer so that they correctly fulfill the query. Ensure the modified code is minimal, clean, and logically sound.
3. Ambiguity and Tie Cases
In instances where the query yields no unique resolution - whether due to tied outcomes, under specified conditions, or semantic ambiguity in the formulation - you can employ principled strategies to ensure meaningful processing:
- Option A: Modify the query to resolve the ambiguity (e.g., make clarification).
- Option B: If the ambiguity is irreparable or the answer depends on arbitrary choices, discard the case.
Additional Guidelines
Maintain consistency between the query, code, and answer. Ensure your corrections do not introduce new ambiguities or assumptions not grounded in the table or query.
B.3 Guidelines for Document Reference
Objective
Your task is to add an additional reasoning hop to the original query using the provided Wikipedia entities and content. This will help transform the original query into a more complex, multi-hop question that requires deeper reasoning.
Input Data
Each data case consists of:
Original Query
Answer
A set of Wikipedia entities relevant to the query or the answer
Corresponding Wikipedia content for each entity
Task Overview
For each data case, check if the original answer is mentioned in the provided Wiki entities:
- If not mentioned or ambiguous (e.g., "Jordan" for both basketball star and sports brand), do not change the query and answer. Just mark the case as “No modification needed”.
- If the answer entity exists in the Wikipedia content, perform the following steps:
1. Identify key factual descriptions about the answer entity from its Wiki content.
2. Add a reasoning hop to the original query that leads to the answer via this key fact.
3. Generate two candidate (Query, Answer) pairs by rewriting the query to incorporate this extra reasoning step and updating the answer accordingly.
Example
Original Query: Which album takes first place on the Billboard leaderboard in 2013?
Original Answer: ArtPop
Wiki Entity: ArtPop (album)
A key description: It was released on November 6, 2013, by Streamline and Interscope Records.
Modified Query Candidates:
Who released the album that takes the first place on the Billboard leaderboard in 2013?
→ Answer: Streamline and Interscope Records
Appendix C Implementation Details of Baselines
This section provides a detailed account of the implementation procedures for the baseline methodologies to ensure a fair and reproducible evaluation.
C.1 ReAct
For the ReAct framework, we preprocess tabular data by converting it into markdown-formatted plain text. To ensure consistency in experimental conditions, we adopt the same chunking and retrieval configurations as those employed in our TableRAG model. We build upon the publicly available ReAct implementation444https://github.com/ysymyth/ReAct. The framework addresses user queries through an iterative reasoning loop consisting of Thought, Action, and Observation steps, culminating in a final answer via the Finish operation. The max iteration is set to 5, the same as TableRAG.
C.2 TableGPT2
We evaluate TableGPT2 using its officially released TableGPT Agent implementation 555https://github.com/tablegpt/tablegpt-agent. As specified in its API documentation, we provide both the document content and the tabular data as inputs to the HumanMessage class, ensuring adherence to the intended usage of the model:
| table | List_of_Australian_films_of_2012_0 |
| query | Who wrote and starred the comedy film released in the second half of 2012 (July-December) that had the highest number of cast members in the List of Australian films of 2012? |
| sql_query | Which comedy film released in the second half of 2012 (July-December) had the highest number of cast members in the List of Australian films of 2012? |
[…]
| sql_ans | Kath & Kimderella |
| answer | Riley, Turner, and Magda Szubanski |
Appendix D Analysis
We present a series of auxiliary analyses conducted on both our proposed benchmark, HeteQA, and the publicly available HybridQA dataset. These analyses offer further insight into the behavior and limitations of TableRAG beyond the primary evaluation metrics.
D.1 Analysis on HybridQA
We extend our investigation to the HybridQA dataset by examining the performance distribution on data domains and conducting a detailed error analysis, using DeepSeek-v3 as the backbone model. The results, summarized in Figures 10 and 9, reveal patterns consistent with those observed on HeteQA. This parallel further validates the generality of our observations across HeteQA settings.
D.2 HyperParameter Analysis
To elucidate the influence of the top- retrieval parameter on the performance of TableRAG, we undertook a systematic sensitivity analysis. While our main experimental setup fixed , we expanded our investigation on the HeteQA dataset, utilizing DeepSeek-V3 as the retrieval backbone, and varied across the set . The resulting performance metrics are illustrated in Figure 11.
Our observations reveal distinct behaviors across benchmarks with respect to the choice of . Notably, HybridQA exhibits superior performance at higher values. This effect is plausibly attributable to the lack of deduplication within its tables, whereby overlapping or similar information from multiple tables or Wikipedia documents contributes constructively during retrieval. In aggregate, setting strikes an effective balance, yielding robust performance across both benchmarks while maintaining computational efficiency. This efficiency gain arises from the direct relationship between the top- retrieval size and the subsequent LLM context length, underscoring the practical importance of this hyperparameter in optimizing the trade-off between accuracy and resource consumption.
Appendix E Check List
Harmful information And Privacy
We propose a new RAG solution to address multi-hop problems related to heterogeneous data, without involving any harmful information or privacy. In addition, we provide a heterogeneous benchmark containing tables and texts sourced from Wikipedia. All of the data is publicly available, contains no personal information, and involves no harmful content.
License and Intend
We provide the license we used here:
Claude 3.5 Sonnet (https://www.anthropic.com/legal/aup)
Qwen2.5-72B-Instruct (https://huggingface.co/Qwen/Qwen2.5-72B-Instruct/blob/main/LICENSE)
TableGPT Agent (Apache License 2.0) and TableGPT2-7B (https://huggingface.co/tablegpt/TableGPT2-7B/blob/main/LICENSE)
Our use of these existing artifacts was consistent with their intended use.
Documentation of the artifacts
We propose a novel Retrieval-Augmented Generation (RAG) framework that integrates traditional document retrieval with structured data querying via SQL, aiming to enhance performance in table-related question answering tasks. The framework comprises two fundamental stages: offline database construction and online interactive reasoning. Compared to conventional document-centric RAG approaches, our architecture offers additional reliable SQL execution results—when table fragments are involved in the retrieval process. This supplementary source mitigates the limitations of generative models in handling structured tabular data.
To facilitate a more comprehensive evaluation of our framework, we further construct a heterogeneous benchmark named HeteQA. This benchmark aims to evaluate the capability to handle multi-hop reasoning tasks across heterogeneous documents. The benchmark instances are initially generated by LLMs, and then rigorously validated by experienced programmers and database administrators to ensure correctness and realism.
Appendix F Prompts
The prompts for TableRAG are presented in this section.

### v2

@@
 The prevailing practice of flattening tables and chunking strategies disrupts the intrinsic tabular structure, leads to information loss, and undermines the reasoning capabilities of LLMs in multi-hop, global queries.
-To address these challenges, we propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data.
+To address these challenges, we propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data.
 TableRAG iteratively operates in four steps: context-sensitive query decomposition, text retrieval, SQL programming and execution, and compositional intermediate answer generation.
@@
 Experimental results demonstrate that TableRAG consistently outperforms existing baselines on both public datasets and our HeteQA, establishing a new state-of-the-art for heterogeneous document question answering.
-We release TableRAG at https://github.com/yxh-y/TableRAG/tree/main.
 TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning
-Xiaohan Yu∗ Huawei Cloud BU, Beijing yuxiaohan5@huawei.com Pu Jian∗ Huawei Cloud BU, Beijing jianpu2@huawei.com Chong Chen† Huawei Cloud BU, Beijing chenchong55@huawei.com
+Xiaohan Yu∗, Pu Jian∗, Chong Chen#
+Huawei Cloud BU, Beijing
+{yuxiaohan5, jinapu2, chenchong55}@huawei.com
+https://github.com/yxh-y/TableRAG
 1 Introduction
@@
 As illustrated in Figure 1, the RAG approach computes percentage over the top-N most relevant chunks rather than the full table, and thus results in an incorrect answer.
-To address these limitations of existing RAG systems, we propose TableRAG, an hybrid (SQL execution and text retrieval) framework that dynamic transitions between textual understanding and complex manipulations over tabular data.
+To address these limitations of existing RAG systems, we propose TableRAG, an SQL-based framework that dynamically transitions between textual understanding and complex manipulations over tabular data.
 TableRAG interacts with tables by leveraging SQL as an interface.
@@
 We identify two key limitations of existing RAG approaches in the context of heterogeneous document question answering: structural information loss and lack of global view.
-We propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data.
+We propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data.
 TableRAG comprises an offline database construction phase and a four-step online iterative reasoning process.
@@
 3.1 Overview and Design Principles
-We propose TableRAG, an hybrid (SQL execution and text retrieval) framework designed to preserve table structural integrity and facilitate heterogeneous reasoning.
+We propose TableRAG, an SQL-based framework designed to preserve table structural integrity and facilitate heterogeneous reasoning.
 As depicted in Figure 2, TableRAG consists of offline and online workflows.
@@
 Through repeated cycles of decomposition and resolution, a solution to the query is progressively constructed.
-Detailed prompt templates are provided in Appendix F.
+Detailed prompt templates are provided in Appendix LABEL:prompts.
 Context-Sensitive Query Decomposition
@@
 This entity substitution can either modify the subject of the question and its corresponding answer or alter the query phrasing while preserving the original answer.
-The annotation guidelines and annotator profiles are detailed in Appendix B.
+The annotation guidelines and annotator profiles are detailed in Appendix LABEL:instructions.
 4.2 Discussion
@@
 We use accuracy as the evaluation metric, assessed by Qwen-2.5-72B, which yields a binary score of 0 or 1.
-The prompt is shown in Appendix F.
+The prompt is provided in Appendix LABEL:prompts, and the results evaluated using the exact match metric are presented in Appendix LABEL:append_exp.
 5.1.3 Baselines
@@
 (4) TableGPT2 (Su et al., 2024) employs a Python-based execution module to generate code (e.g., Pandas) for answer derivation within a simulated environment.
-The detailed implementation of these baseline methods is provided in Appendix C.
+The detailed implementation of these baseline methods is provided in Appendix LABEL:baseline and additional baseline comparisons are presented in Appendix LABEL:append_exp.
 5.2 Main Result
@@
 This can be attributed to context insensitivity during multi-turn reasoning.
-Queries decomposed into multiple sub-tasks, such as filtering or aggregation, suffering from information incompleteness or error propagation.
+Queries that could be resolved with single SQL execution are instead decomposed into multiple subqueries.
+This over-fragmentation may introduce errors or incomplete information during execution (especially when operations such as filtering, aggregation are involved), potentially leading to cascading failures.
 (2) TableGPT2 yields acceptable results solely on single-source queries, such as WikiTQ, underscoring its limited capacity for handling multi-source queries.
@@
 Moreover, the consistency in performance across different LLM backbones underscores TableRAG’s architectural generality and compatibility with a broad range of backbones.
+| Method | WikiTQ | HeteQA | ||
+| Total Latency | Avg. Latency / Step | Total Latency | Avg. Latency / Step | |
+| ReAct | 13.50 | 5.92 | 24.57 | 7.70 |
+| TableRAG | 17.70 | 7.93 | 45.65 | 10.79 |
 5.3 Ablation Study
@@
 6 Efficiency
+6.1 Execution Dynamics
 We evaluate the efficiency of TableRAG by examining the distribution of its execution iterations, as illustrated in Figure 5.
@@
 It is attributed to the incorporation of SQL-based tabular reasoning.
+6.2 Latency
+We assess the latency performance using Qwen-72B-Instruct as a consistent backbone model.
+As shown in Table 2, TableRAG incurs lower latency compared to ReAct, in both average time per step and total time efficiency.
+This is mainly due to: (1) TableRAG executes fewer iterations, and employs symbolic SQL execution, which is computationally efficient.
+(2) ReAct performs reasoning at every step, and consumes more steps before reaching a final decision, thereby incurring greater time overhead.
 7 Analysis
-We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix D.
+We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix LABEL:analysis.
 7.1 Error Analysis
@@
 Appendix A HeteQA
-A.1 Prompt Design
-The templates employed for HeteQA benchmark construction are presented in Table LABEL:SQL.
-A.2 Table Collection
-To enable complex reasoning over tabular structures, we curate a subset of extensive tables from the Wikipedia-based corpus.
-Specifically, we retain only tables containing more than 20 rows and at least 7 columns.
-This filtering process reduced the initial collection from 15,314 tables to 1,345 candidates that meet the criteria for structural richness.
-To further enhance query diversity and reduce redundancy in the dataset, we apply a deduplication step based on schema similarity.
-In cases where multiple tables share identical column structures — for example, the entries for 1947 BAA draft and 1949 BAA draft — only a single representative instance is preserved.
-This procedure yielded a final dataset comprising 155 unique tables.
-A.3 Data Collection
-As illustrated in Figure 8, each data instance is constructed through a three-stage pipeline, facilitated via the Claude 3.7 Sonnet API333https://www.anthropic.com/claude/sonnet.
-To construct the dataset, we prompt the LLM to generate SQL or pandas code associated with queries conditioned on a provided table schema.
-The generated code is then executed against the table to obtain the corresponding answer.
-Instances that fail to produce an executable result are discarded.
-To promote diversity within the dataset and avoid redundancy, we eliminate any instances exhibiting duplicate queries or identical answers.
-This filtering is performed using regular expressions to detect lexical or semantic repetition.
-Additionally, we discard cases where the execution result is ambiguous, such as queries seeking the top-ranked entity when multiple candidates are tied.
-Following automated filtering, all remaining instances undergo a manual verification process.
-Two human annotators independently assess each query, the corresponding code, and the resulting answer.
-In cases of inconsistency or error, annotators revise the SQL or Python code and correct the associated answer to ensure accuracy and coherence.
-Appendix B Guidelines for HeteQA Annotations
-B.1 Annotator Profiles
-The two annotators are volunteers whose native language is Chinese and who possess fluent English proficiency.
-Both annotators are professional software engineers with substantial experience in SQL and Python programming.
-B.2 Guidelines for Answer Verification
-Objective
-Your task is to verify whether the provided answer correctly and completely addresses the natural language query, based on the associated code and table.
-You will also identify and handle ambiguous or non-unique cases.
-Annotation Steps
-1. Verify the Answer
-Read the natural language query and examine the accompanying Excel table, SQL statements, and Python (pandas) code.
-Verify whether the output aligns with the semantics of the original query and the underlying data.
-To this end, you are permitted to execute multiple operations on the provided Excel files.
-2. Correction Tasks (if necessary)
-If the answer is incorrect, modify both the code and the answer so that they correctly fulfill the query. Ensure the modified code is minimal, clean, and logically sound.
-3. Ambiguity and Tie Cases
-In instances where the query yields no unique resolution - whether due to tied outcomes, under specified conditions, or semantic ambiguity in the formulation - you can employ principled strategies to ensure meaningful processing:
-- Option A: Modify the query to resolve the ambiguity (e.g., make clarification).
-- Option B: If the ambiguity is irreparable or the answer depends on arbitrary choices, discard the case.
-Additional Guidelines
-Maintain consistency between the query, code, and answer. Ensure your corrections do not introduce new ambiguities or assumptions not grounded in the table or query.
-B.3 Guidelines for Document Reference
-Objective
-Your task is to add an additional reasoning hop to the original query using the provided Wikipedia entities and content.
-This will help transform the original query into a more complex, multi-hop question that requires deeper reasoning.
-Input Data
-Each data case consists of:
-Original Query
-Answer
-A set of Wikipedia entities relevant to the query or the answer
-Corresponding Wikipedia content for each entity
-Task Overview
-For each data case, check if the original answer is mentioned in the provided Wiki entities:
-- If not mentioned or ambiguous (e.g., "Jordan" for both basketball star and sports brand), do not change the query and answer. Just mark the case as “No modification needed”.
-- If the answer entity exists in the Wikipedia content, perform the following steps:
-1. Identify key factual descriptions about the answer entity from its Wiki content.
-2. Add a reasoning hop to the original query that leads to the answer via this key fact.
-3. Generate two candidate (Query, Answer) pairs by rewriting the query to incorporate this extra reasoning step and updating the answer accordingly.
-Example
-Original Query: Which album takes first place on the Billboard leaderboard in 2013?
-Original Answer: ArtPop
-Wiki Entity: ArtPop (album)
-A key description: It was released on November 6, 2013, by Streamline and Interscope Records.
-Modified Query Candidates:
-Who released the album that takes the first place on the Billboard leaderboard in 2013?
-→ Answer: Streamline and Interscope Records
-Appendix C Implementation Details of Baselines
-This section provides a detailed account of the implementation procedures for the baseline methodologies to ensure a fair and reproducible evaluation.
-C.1 ReAct
-For the ReAct framework, we preprocess tabular data by converting it into markdown-formatted plain text.
-To ensure consistency in experimental conditions, we adopt the same chunking and retrieval configurations as those employed in our TableRAG model.
-We build upon the publicly available ReAct implementation444https://github.com/ysymyth/ReAct.
-The framework addresses user queries through an iterative reasoning loop consisting of Thought, Action, and Observation steps, culminating in a final answer via the Finish operation.
-The max iteration is set to 5, the same as TableRAG.
-C.2 TableGPT2
-We evaluate TableGPT2 using its officially released TableGPT Agent implementation 555https://github.com/tablegpt/tablegpt-agent.
-As specified in its API documentation, we provide both the document content and the tabular data as inputs to the HumanMessage class, ensuring adherence to the intended usage of the model:
 | table | List_of_Australian_films_of_2012_0 |
@@
 | sql_query | Which comedy film released in the second half of 2012 (July-December) had the highest number of cast members in the List of Australian films of 2012? |
-[…]
-| sql_ans | Kath & Kimderella |
-| answer | Riley, Turner, and Magda Szubanski |
-Appendix D Analysis
-We present a series of auxiliary analyses conducted on both our proposed benchmark, HeteQA, and the publicly available HybridQA dataset.
-These analyses offer further insight into the behavior and limitations of TableRAG beyond the primary evaluation metrics.
-D.1 Analysis on HybridQA
-We extend our investigation to the HybridQA dataset by examining the performance distribution on data domains and conducting a detailed error analysis, using DeepSeek-v3 as the backbone model.
-The results, summarized in Figures 10 and 9, reveal patterns consistent with those observed on HeteQA.
-This parallel further validates the generality of our observations across HeteQA settings.
-D.2 HyperParameter Analysis
-To elucidate the influence of the top- retrieval parameter on the performance of TableRAG, we undertook a systematic sensitivity analysis.
-While our main experimental setup fixed , we expanded our investigation on the HeteQA dataset, utilizing DeepSeek-V3 as the retrieval backbone, and varied across the set .
-The resulting performance metrics are illustrated in Figure 11.
-Our observations reveal distinct behaviors across benchmarks with respect to the choice of .
-Notably, HybridQA exhibits superior performance at higher values.
-This effect is plausibly attributable to the lack of deduplication within its tables, whereby overlapping or similar information from multiple tables or Wikipedia documents contributes constructively during retrieval.
-In aggregate, setting strikes an effective balance, yielding robust performance across both benchmarks while maintaining computational efficiency.
-This efficiency gain arises from the direct relationship between the top- retrieval size and the subsequent LLM context length, underscoring the practical importance of this hyperparameter in optimizing the trade-off between accuracy and resource consumption.
-Appendix E Check List
-Harmful information And Privacy
-We propose a new RAG solution to address multi-hop problems related to heterogeneous data, without involving any harmful information or privacy.
-In addition, we provide a heterogeneous benchmark containing tables and texts sourced from Wikipedia.
-All of the data is publicly available, contains no personal information, and involves no harmful content.
-License and Intend
-We provide the license we used here:
-Claude 3.5 Sonnet (https://www.anthropic.com/legal/aup)
-Qwen2.5-72B-Instruct (https://huggingface.co/Qwen/Qwen2.5-72B-Instruct/blob/main/LICENSE)
-TableGPT Agent (Apache License 2.0) and TableGPT2-7B (https://huggingface.co/tablegpt/TableGPT2-7B/blob/main/LICENSE)
-Our use of these existing artifacts was consistent with their intended use.
-Documentation of the artifacts
-We propose a novel Retrieval-Augmented Generation (RAG) framework that integrates traditional document retrieval with structured data querying via SQL, aiming to enhance performance in table-related question answering tasks.
-The framework comprises two fundamental stages: offline database construction and online interactive reasoning.
-Compared to conventional document-centric RAG approaches, our architecture offers additional reliable SQL execution results—when table fragments are involved in the retrieval process.
-This supplementary source mitigates the limitations of generative models in handling structured tabular data.
-To facilitate a more comprehensive evaluation of our framework, we further construct a heterogeneous benchmark named HeteQA.
-This benchmark aims to evaluate the capability to handle multi-hop reasoning tasks across heterogeneous documents.
-The benchmark instances are initially generated by LLMs, and then rigorously validated by experienced programmers and database administrators to ensure correctness and realism.
-Appendix F Prompts
-The prompts for TableRAG are presented in this section.
+| sql |